In a quick return motion mechanism of a Crank and slotted Lever type the ratio of maximum velocities is 2

1
Expert's answer
2022-02-23T13:16:03-0500

"(a)\ \dfrac{x}{25} = 2"


"x = 50\ cm"


"(b)\ \sin\dfrac{25}{50}= \sin({90-\dfrac{\theta}2)}"


"30° = 90 -\dfrac{\theta}2"


"\theta= 120°"


ratio of time of cutting to return strokes = "\dfrac{180-\theta}{\theta}"


"(c)\ 30\ rpm × 2 = 60\ rpm"